OpenStudy (anonymous):

Which is the equation of a circle with center (–2, 3) and radius r = 3

OpenStudy (smored):

(x-a)^2 + (y-b)^2 = r2 a=x-coord for center b=y-coord for center r=radius

OpenStudy (anonymous):

^ Plug your information into that general equation and you're good to go.

OpenStudy (smored):

\[(x-a)^2 + (y-b)^2 = r^2\]

OpenStudy (anonymous):

I don't know how to do this may you work it out?

jimthompson5910 (jim_thompson5910):